OpenStudy (anonymous):

∫θcosθdΘu=θdv=cosΘdΘdu=dΘv=sinΘ I got: ∫ΘcosThetadΘ=sin2Θ+cosΘ+C Is this correct? Thanks:)

OpenStudy (turingtest):

\[\int\theta\cos\theta d\theta\]is the problem?

OpenStudy (anonymous):

yes. u=theta dv=costheta dtheta

OpenStudy (turingtest):

I'm don't think that;'s quite right, but let me do it\[u=\theta\implies du=d\theta\]\[dv=\cos\theta d\theta\implies v=\sin\theta\]\[\int\theta\cos\theta=\theta\sin\theta-\int\sin\theta d\theta=\theta\sin\theta+\cos\theta\]

OpenStudy (turingtest):

...+C

OpenStudy (turingtest):

do you follow? do you see your mistake?

OpenStudy (anonymous):

But \[\Theta(\sin \Theta)=\sin^2\Theta\] and the integral of sin=(-cos)??

OpenStudy (turingtest):

\[\theta\sin\theta\neq\sin^2\theta\]

OpenStudy (turingtest):

and I applied the fact that integral of sin is -cos by changing the - to a + between the two integrals in my final answer

OpenStudy (anonymous):

ok, so I just did my algebra wrong by combining the sin theta and theta.

OpenStudy (turingtest):

yeah, that's a no-no...

OpenStudy (turingtest):

there is no identity I know of for\[x\sin x\]

OpenStudy (anonymous):

Got you, Appreciate your help. I'm trying to study for Calc2 in Aug, lol.

OpenStudy (turingtest):

Great, good luck. I hope you come to Open Study for help :)